How many atoms are in 1.8675g of a recently synthesized organic compound:

1.which has a molar mass between 350 and 400 grams/mole
2.which consists of only iron, carbon and hydrogen
3.when analyzed using combustion analysis on a 0.9437 g sample, produced 2.233 g of CO2 and 0.457 g of H2O?
1
Expert's answer
2012-10-08T09:58:02-0400
We have compound with& formula (C)x(H)y(Fe)o
Let's calculate x,y and o using combustion analysis

2.233 & m=0.6090 g n=0.6090/12= 0.0508
CO2 --------& C
& 44 12

0.457 m=0.0508 g n=0.0508/1=0.0508 mol
H2O------------2H
18 2*1


m of Fe =0.9437-0.609-0.0508=0.2839 n=0.2839/56= 0.00510

n of C : n of H : n of Fe =10:10:1,& it is something like this (C10H10Fe)z& together it have molar mass& (120+10+56)z= 186z

z=400/186=2 (it's closer to 2) , so it is C20H20Fe2

& 1.8672 & N
C20H20Fe2------------- 42 atoms
& 372 42*6.02 x 10^23


N=1.8672* 42*6.02 x 10^23 /& 372 =3.02 x 10^21 atoms